OpenStudy (anonymous):

For the function f(x)=√( x)(x-3) , find the open intervals on which the function is increasing or decreasing, and find the relative extrema. I have tried all that my notes can provide and OS is my last resort. If someone could show me a step by step solution on how to get this done I'd appreciate it so muchh! I really do not know how to continue on with this or if I am in the right direction!

OpenStudy (anonymous):

is the function \[f(x)=\sqrt{x}(x-3)\]?

OpenStudy (anonymous):

or is it all under the radical like \[f(x)=\sqrt{x(x-3)}\]?

OpenStudy (anonymous):

the first one :)

OpenStudy (anonymous):

Do I start with finding the 2nd derivative? I just feel clueless at this point.

OpenStudy (anonymous):

\[f'(x)=\sqrt{x}.1+\frac{ x-3 }{2\sqrt{x} }=\frac{ 2x+x-3 }{2\sqrt{x} }=\frac{ 3x-3 }{ 2 \sqrt{x} }\] f'(x)=0 gives x=1 \[f \prime \prime \left( x \right)=\frac{ 1 }{2 } *\frac{\sqrt{x}*3-(3x-3)\frac{ 1 }{2 \sqrt{x}} }{x }\] at x=1 \[f \prime \prime(x)=\frac{ 1 }{2}*\frac{ \sqrt{1}*3-0 }{ 2*\sqrt{1} }=\frac{ 3 }{4 }>0\] Hence f(x) is minimum at x=1 for increasing f'(x)>0 for decreasing f'(x)<0 solve.

OpenStudy (anonymous):

remember here \[x \ge 0\]